OpenStudy (anonymous):

how can I share five(5) oranges between two people

OpenStudy (anonymous):

equally?

OpenStudy (anonymous):

If you're able to cut those oranges up, infinitely many ways.

OpenStudy (lgbasallote):

well you can say there are just four...if you know what i mean ;D

OpenStudy (diyadiya):

1+4=5 2+3=5 =D you gotta share equally?

OpenStudy (lgbasallote):

you forgot your favorite kind of sharing @Diyadiya 0 + 5 = 5 :DDD

OpenStudy (diyadiya):

@lgbasallote lol you SHARE that way ?

OpenStudy (lgbasallote):

hahaha don't we all :P

OpenStudy (anonymous):

I can't eat 2 and a half oranges, so you can have 4. There! Easy!